# Maximum entropy distributions of wavefunctions at thermal equilibrium

### 19 Mar 2026 v1

@@ Conclusion
$\frac{\int[d\Gamma]\exp\left(-\beta^{\prime}\mathrm{Tr}\{H\Gamma\}\right)}{\int[d\Gamma]}$ $=\int_{U(N)}[dU]\exp\left(-\beta^{\prime}\mathrm{Tr}\{HU\Gamma_{0}U^{\dagger}\}\right),$ (S9)
@@
$P_{\mathrm{GCE}}(\Gamma)=\frac{\exp\left(-\sum_{n=1}^{N}\mu_{n}\Gamma_{nn}-\sum_{n<m}\eta^{\mathrm{R}}_{mn}\Gamma_{mn}^{\mathrm{R}}-\sum_{n<m}\eta^{\mathrm{I}}_{mn}\Gamma_{mn}^{\mathrm{I}}\right)}{\int[d\Gamma]\exp\left(-\sum_{n=1}^{N}\mu_{n}\Gamma_{nn}-\sum_{n<m}\eta^{\mathrm{R}}_{mn}\Gamma_{mn}^{\mathrm{R}}-\sum_{n<m}\eta^{\mathrm{I}}_{mn}\Gamma_{mn}^{\mathrm{I}}\right)},$ (S20)
@@
$\mu_{i}\approx-F(\beta)p_{i}+\frac{1}{p_{i}},\quad\mathrm{where}\ \lim_{\beta\rightarrow 0}F(\beta)=N,\ \mathrm{and}\ \lim_{\beta\rightarrow\infty}F(\beta)=1,$ (S36)
@@
where $C(\rho)$ is the constraint function, and $D_{\alpha}(\Gamma\|\rho)$ is the $\alpha$-Rényi divergence of wavefunction $\Gamma$ from the density operator $\rho$. Substituting eq.˜S38 into eq.˜S4 yields the maximum entropy distribution,
@@
$=\frac{1}{\Gamma(\mu)}\sum_{i}\frac{1}{\prod_{j\neq i}(\lambda_{j}^{-1}-\lambda_{i}^{-1})}\int_{0}^{\infty}dt\ t^{\mu-N}e^{-t\lambda_{i}^{-1}},$ (S44)
@@
$C(\rho)=Q(\rho)+H_{N}-1,\quad Q(\rho)=-\sum_{k}\frac{p_{k}^{N}\ln p_{k}}{\prod_{i\neq k}(p_{k}-p_{i})},\quad H_{N}=\sum_{k=1}^{N}\frac{1}{k}.$ (S56)
@@
that is consistent with the Gibbs state (satisfying $\langle\Gamma\rangle_{P}=\rho,\ \forall\rho$). In other words, self-consistency requires that the density operator which we constrain the Rényi divergence on, to be equal to the density operator of the corresponding maximum entropy distribution. In the main text, we set $\alpha=2$ and $\mu=N+1$ to coincide with the Scrooge ensemble, without any a priori principle to select this constraint; however, the following uniqueness property proves that there is no other self-consistent alternative. The population of the $i$-th state in the distribution $P_{\alpha\mu}(\Gamma)$ is given by (proof in section S3.3 of the SI),
$\tilde{p}_{k}(\alpha,\mu):=\langle\Gamma_{kk}\rangle_{P_{\alpha\mu}}=\frac{\lambda_{k\alpha}}{\mu-N}\frac{\partial}{\partial\lambda_{k\alpha}}\ln\sum_{i}\frac{\lambda_{i\alpha}^{\mu-1}}{\prod_{j\neq i}(\lambda_{i\alpha}-\lambda_{j\alpha})}$ (S58)
where $\lambda_{k\alpha}=p_{k}^{\alpha-1}$, and we assume that $\alpha\neq 1$. A parameter set $(\alpha,\mu)$ is consistent with the Gibbs state if the density operator corresponding to $P_{\alpha\mu}$ is equal to the density operator $\rho$ in the constraint (i.e., $\tilde{p}_{k}(\alpha,\mu)=p_{k},\ \forall k,\rho$). Assuming that $(\alpha,\mu)$ is consistent with the Gibbs state, we can integrate the $N$ partial derivatives eq.˜S58 to yield,
@@
where $c$ is a constant that is independent of $\rho$. Since the left-hand side of eq.˜S59 is independent of $\rho$, the right-hand side must also be independent. Using identities from Lagrange interpolants, we have that
@@
The case $\mu=N+1$ is independent of $\rho$ only when $\alpha=2$, such that $\sum_{i}\lambda_{i\alpha}=\sum_{i}p_{i}=1$. The cases $\mu=\{0,1,\dots,N\}$ correspond to spurious solutions from the differentiation and integration of eq.˜S58, as can be confirmed by numerical solutions. Figure˜S3 shows that the variance of the $N$ Lagrange multipliers obtained by solving the populations $\tilde{p}_{k}(\alpha,\mu_{k})=p_{k}$ only vanishes at $\alpha=2$, where $\mu_{k}=N+1,\ \forall k$.
@@
Figure S3: There is a unique Rényi divergence constraint $\langle D_{\alpha}(\Gamma\|\rho)\rangle_{P}=C(\rho)$ where the maximum entropy distribution has an average $\tilde{\rho}$ that is equal to $\rho$. (a) The NN Lagrange multipliers $\{\mu_{k}\}$ obtained by numerically solving the $N$ population constraints $\tilde{\rho}=\rho$ (eq.˜S58) as a function of $\alpha$. A valid solution requires that $\{\mu_{k}\}$ be identical, which is only satisfied at $\alpha=2$, where $\mu_{k}=N+1,\ \forall k$. (b) The variance of $\{\mu_{k}\}$ only vanishes at $\alpha=2$, for all density matrices and temperatures $\beta$. Results are calculated for a Hamiltonian spectrum $\sigma\{H\}=\{0,1,2\}$.
@@
$p_{k}^{\mathrm{ECE}}$ $=-\frac{\partial}{\partial(\beta^{\prime}E_{k})}\ln\left(\sum_{i=1}^{N}\alpha_{i}\right),\quad\alpha_{i}=\frac{e^{-\beta^{\prime}E_{i}}}{\prod_{j\neq i}(\beta^{\prime}E_{j}-\beta^{\prime}E_{i})},$ (S65)
@@
$=\frac{(2\pi)^{N-1}}{Z_{\alpha\mu}}\frac{1}{\Gamma(\mu)}\int_{0}^{\infty}dt\ t^{\mu-1}\int_{\Delta^{N-1}}dx\ x_{k}\ e^{-t\sum_{i}x_{i}\lambda_{i\alpha}^{-1}},$ (S71)
$=\frac{(2\pi)^{N-1}}{Z_{\alpha\mu}}\frac{1}{\Gamma(\mu)}\int_{0}^{\infty}dt\ t^{\mu-N-1}\left[-\frac{\partial}{\partial\lambda_{k\alpha}^{-1}}\sum_{i}\frac{e^{-t\lambda_{i\alpha}^{-1}}}{\prod_{j\neq i}(\lambda_{j\alpha}^{-1}-\lambda_{i\alpha}^{-1})}\right],$ (S72)
$=\frac{\displaystyle\left(\mu-1+\sum_{i\neq k}\frac{\lambda_{k\alpha}}{\lambda_{i\alpha}-\lambda_{k\alpha}}\right)w_{k}(\alpha,\mu)+\sum_{i\neq k}w_{i}(\alpha,\mu)\frac{\lambda_{k\alpha}}{\lambda_{i\alpha}-\lambda_{k\alpha}}}{\displaystyle(\mu-N)\sum_{i}w_{i}(\alpha,\mu)},$ (S73)
@@
where we have used eq.˜S47, eq.˜S48, eq.˜S49, and defined
@@
$S_{\mathrm{ECE}}=\beta^{\prime}U+\ln\mathcal{Z}_{\mathrm{ECE}}$ (S77)
$S_{\mathrm{GCE}}=\mathrm{Tr}\{M\rho\}+\ln\mathcal{Z}_{\mathrm{GCE}}$ (S78)
@@
$\mathcal{Z}_{\mathrm{Scr}}=\int[d\Gamma](\mathrm{Tr}\{\Gamma\rho^{-1}\})^{-(N+1)}=\frac{(2\pi)^{N-1}\det\rho}{N!}.$ (S80)
@@
$\lim_{\beta\rightarrow\infty}(S_{\mathrm{GCE}}-S_{\mathrm{Scr}})$ $=\lim_{\beta\rightarrow\infty}\left[\ln\left(\frac{\mathcal{Z}_{\mathrm{GCE}}}{\mathcal{Z}_{\mathrm{Scr}}}\right)+\mathrm{Tr}\{M\rho\}-(N+1)(Q(\rho)+H_{N}-1)\right]$ (S81)
$=\ln\left(\lim_{\beta\rightarrow\infty}\frac{\mathcal{Z}_{\mathrm{GCE}}}{\mathcal{Z}_{\mathrm{Scr}}}\right)+(N-1)-(N+1)(H_{N}-1)$ (S82)
@@
where $S_{\mathrm{Haar}}=\ln\mathcal{Z}_{\mathrm{Haar}}$, and $\mathcal{Z}_{\mathrm{Haar}}=\int[d\Gamma]=\frac{(2\pi)^{N-1}}{(N-1)!}$. Equation˜S89 is the entropy difference incurred by a $\rho$-distortion i.e., the method used by Jozsa et al. to derive the Scrooge ensemble from deforming the Haar ensemble.[27]
Figure S4: The Scrooge ensemble only converges to the GCE in the infinite temperature limit. (a) The ensemble entropy of the GCE ($S_{\mathrm{GCE}}$) is always greater than or equal to the entropy of the Scrooge ensemble ($S_{\mathrm{Scr}}$). (b) The entropy difference $\Delta S=S_{\mathrm{GCE}}-S_{\mathrm{Scr}}$ approaches the limit eq.˜S87 (dashed lines) at low temperatures. The difference $\Delta S$ increases with $N$, showing that convergence is not achieved in the thermodynamic limit. Results are calculated for Hamiltonian spectra $\sigma\{H\}=\{k\}_{k=0}^{N-1}$.
@@
Figure S5: The ECE exhibits ground state condensation in the thermodynamic limit. (a) The energy level diagram of an NN-level system with two low-energy states at 0 and $\delta$, and a set of high-energy levels above $\Delta\gg\delta$. (b) The ground state population in the ECE (red stars) increases with $N$, even though it should be unaffected by the inclusion of thermally-inaccessible levels, like the Gibbs state (black). Populations agree with Monte Carlo sampling (pink circles). Results are calculated for $\delta=0.1$, $\Delta=100$, and $\beta=1$.

### 20 Mar 2026 v2

@@ Conclusion
$\frac{\int[d\Gamma]e^{-\beta^{\prime}\mathrm{Tr}\{H\Gamma\}}}{\int[d\Gamma]}$ $=\int_{U(N)}[dU]\exp\left(-\beta^{\prime}\mathrm{Tr}\{HU\Gamma_{0}U^{\dagger}\}\right),$ (S9)
@@
$P_{\mathrm{GCE}}(\Gamma)=\frac{e^{-\sum_{n=1}^{N}\mu_{n}\Gamma_{nn}-\sum_{n<m}\eta^{\mathrm{R}}_{mn}\Gamma_{mn}^{\mathrm{R}}-\sum_{n<m}\eta^{\mathrm{I}}_{mn}\Gamma_{mn}^{\mathrm{I}}}}{\int[d\Gamma]e^{-\sum_{n=1}^{N}\mu_{n}\Gamma_{nn}-\sum_{n<m}\eta^{\mathrm{R}}_{mn}\Gamma_{mn}^{\mathrm{R}}-\sum_{n<m}\eta^{\mathrm{I}}_{mn}\Gamma_{mn}^{\mathrm{I}}}},$ (S20)
@@
$\mu_{i}\approx-F(\beta)p_{i}+\frac{1}{p_{i}},$ (S36)
where
$\lim_{\beta\rightarrow 0}F(\beta)=N,\ \mathrm{and}\ \lim_{\beta\rightarrow\infty}F(\beta)=1,$ (S37)
@@
where $C(\rho)$ is the constraint function, and $D_{\alpha}(\Gamma\|\rho)$ is the $\alpha$-Rényi divergence of wavefunction $\Gamma$ from the density operator $\rho$. Substituting eq.˜S39 into eq.˜S4 yields the maximum entropy distribution,
@@
$=\frac{1}{\Gamma(\mu)}\sum_{i}\frac{\int_{0}^{\infty}dt\ t^{\mu-N}e^{-t\lambda_{i}^{-1}}}{\prod_{j\neq i}(\lambda_{j}^{-1}-\lambda_{i}^{-1})},$ (S45)
@@
$C(\rho)=Q(\rho)+H_{N}-1$ (S57)
where
$Q(\rho)=-\sum_{k}\frac{p_{k}^{N}\ln p_{k}}{\prod_{i\neq k}(p_{k}-p_{i})}$ (S58)
is the subentropy, and $H_{N}=\sum_{k=1}^{N}k^{-1}$ is the $N$-th harmonic number.
@@
that is consistent with the Gibbs state (satisfying $\langle\Gamma\rangle_{P}=\rho,\ \forall\rho$). In other words, self-consistency requires that the density operator which we constrain the Rényi divergence on, to be equal to the density operator of the corresponding maximum entropy distribution. In the main text, we set $\alpha=2$ and $\mu=N+1$ to coincide with the Scrooge ensemble, without any a priori principle to select this constraint; however, the following uniqueness property proves that there is no other self-consistent alternative. The population of the $k$-th state in the distribution $P_{\alpha\mu}(\Gamma)$ is given by (proof in section S3.3 of the SI),
$\tilde{p}_{k}(\alpha,\mu)$ $:=\langle\Gamma_{kk}\rangle_{P_{\alpha\mu}}$ (S60)
$=\frac{\lambda_{k\alpha}}{\mu-N}\frac{\partial}{\partial\lambda_{k\alpha}}\ln\sum_{i}\frac{\lambda_{i\alpha}^{\mu-1}}{\prod_{j\neq i}(\lambda_{i\alpha}-\lambda_{j\alpha})}$ (S61)
where $\lambda_{k\alpha}=p_{k}^{\alpha-1}$, and we assume that $\alpha\neq 1$. A parameter set $(\alpha,\mu)$ is consistent with the Gibbs state if the density operator corresponding to $P_{\alpha\mu}$ is equal to the density operator $\rho$ in the constraint (i.e., $\tilde{p}_{k}(\alpha,\mu)=p_{k},\ \forall k,\rho$). Assuming that $(\alpha,\mu)$ is consistent with the Gibbs state, we can integrate the $N$ partial derivatives eq.˜S61 to yield,
@@
where $c$ is a constant that is independent of $\rho$. Since the left-hand side of eq.˜S62 is independent of $\rho$, the right-hand side must also be independent. Using identities from Lagrange interpolants, we have that
@@
The case $\mu=N+1$ is independent of $\rho$ only when $\alpha=2$, such that $\sum_{i}\lambda_{i\alpha}=\sum_{i}p_{i}=1$. The cases $\mu=\{0,1,\dots,N\}$ correspond to spurious solutions from the differentiation and integration of eq.˜S61, as can be confirmed by numerical solutions. Figure˜S3 shows that the variance of the $N$ Lagrange multipliers obtained by solving the populations $\tilde{p}_{k}(\alpha,\mu_{k})=p_{k}$ only vanishes at $\alpha=2$, where $\mu_{k}=N+1,\ \forall k$.
@@
Figure S3: There is a unique Rényi divergence constraint $\langle D_{\alpha}(\Gamma\|\rho)\rangle_{P}=C(\rho)$ where the maximum entropy distribution has an average $\tilde{\rho}$ that is equal to $\rho$. (a) The NN Lagrange multipliers $\{\mu_{k}\}$ obtained by numerically solving the $N$ population constraints $\tilde{\rho}=\rho$ (eq.˜S61) as a function of $\alpha$. A valid solution requires that $\{\mu_{k}\}$ be identical, which is only satisfied at $\alpha=2$, where $\mu_{k}=N+1,\ \forall k$. (b) The variance of $\{\mu_{k}\}$ only vanishes at $\alpha=2$, for all density matrices and temperatures $\beta$. Results are calculated for a Hamiltonian spectrum $\sigma\{H\}=\{0,1,2\}$.
@@
$p_{k}^{\mathrm{ECE}}$ $=-\frac{\partial}{\partial(\beta^{\prime}E_{k})}\ln\left(\sum_{i=1}^{N}\alpha_{i}\right),$ (S68)
@@
where
$\alpha_{i}=\frac{e^{-\beta^{\prime}E_{i}}}{\prod_{j\neq i}(\beta^{\prime}E_{j}-\beta^{\prime}E_{i})},$ (S70)
@@
$=\frac{(2\pi)^{N-1}}{Z_{\alpha\mu}\Gamma(\mu)}\int_{0}^{\infty}dt\ t^{\mu-1}\int_{\Delta^{N-1}}dx\ x_{k}\ e^{-t\sum_{i}x_{i}\lambda_{i\alpha}^{-1}},$ (S75)
$=\frac{\displaystyle\left(\mu-1+\sum_{i\neq k}\frac{\lambda_{k\alpha}}{\lambda_{i\alpha}-\lambda_{k\alpha}}\right)w_{k}+\sum_{i\neq k}w_{i}\frac{\lambda_{k\alpha}}{\lambda_{i\alpha}-\lambda_{k\alpha}}}{\displaystyle(\mu-N)\sum_{i}w_{i}},$ (S76)
@@
where we have used eq.˜S48, eq.˜S49, eq.˜S50, and defined
@@
Figure S4: The Scrooge ensemble only converges to the GCE in the infinite temperature limit. (a) The ensemble entropy of the GCE ($S_{\mathrm{GCE}}$) is always greater than or equal to the entropy of the Scrooge ensemble ($S_{\mathrm{Scr}}$). (b) The entropy difference $\Delta S=S_{\mathrm{GCE}}-S_{\mathrm{Scr}}$ approaches the limit eq.˜S91 (dashed lines) at low temperatures. The difference $\Delta S$ increases with $N$, showing that convergence is not achieved in the thermodynamic limit. Results are calculated for Hamiltonian spectra $\sigma\{H\}=\{k\}_{k=0}^{N-1}$.
@@
$S_{\mathrm{ECE}}=\beta^{\prime}U+\ln\mathcal{Z}_{\mathrm{ECE}},$ (S80)
$S_{\mathrm{GCE}}=\mathrm{Tr}\{M\rho\}+\ln\mathcal{Z}_{\mathrm{GCE}},$ (S81)
@@
$\mathcal{Z}_{\mathrm{Scr}}$ $=\int[d\Gamma](\mathrm{Tr}\{\Gamma\rho^{-1}\})^{-(N+1)},$ (S83)
$=\frac{(2\pi)^{N-1}\det\rho}{N!}.$ (S84)
@@
$\lim_{\beta\rightarrow\infty}(S_{\mathrm{GCE}}-S_{\mathrm{Scr}})\\ =\lim_{\beta\rightarrow\infty}\Big[\ln\left(\frac{\mathcal{Z}_{\mathrm{GCE}}}{\mathcal{Z}_{\mathrm{Scr}}}\right)+\mathrm{Tr}\{M\rho\}\\ -(N+1)(Q(\rho)+H_{N}-1)\Big]$ (S85)
$\hskip 42.67912pt=\ln\left(\lim_{\beta\rightarrow\infty}\frac{\mathcal{Z}_{\mathrm{GCE}}}{\mathcal{Z}_{\mathrm{Scr}}}\right)+(N-1)\\ -(N+1)(H_{N}-1)$ (S86)
@@
where $S_{\mathrm{Haar}}=\ln\mathcal{Z}_{\mathrm{Haar}}$, and $\mathcal{Z}_{\mathrm{Haar}}=\int[d\Gamma]=\frac{(2\pi)^{N-1}}{(N-1)!}$. Equation˜S93 is the entropy difference incurred by a $\rho$-distortion i.e., the method used by Jozsa et al. to derive the Scrooge ensemble from deforming the Haar ensemble.[27]
@@
Figure S5: The ECE exhibits ground state condensation in the thermodynamic limit. (a) The energy level diagram of an NN-level system with two low-energy states at 0 and $\delta$, and a set of high-energy levels above $\Delta\gg\delta$. (b) The ground state population in the ECE (red stars) increases with $N$, even though it should be unaffected by the inclusion of thermally-inaccessible levels, like the Gibbs state (black). Results are calculated for $\delta=0.1$, $\Delta=100$, and $\beta=1$.